OpenStudy (anonymous):

please help with this problem, please see drawing its easier to read, thank you

OpenStudy (anonymous):

Wheres the drawing?

OpenStudy (anonymous):



OpenStudy (unklerhaukus):

\[\left(x^m\right)^n=x^{m\times n}\] \[\left(yx^m\right)^n=y^mx^{m\times n}\]

OpenStudy (unklerhaukus):

\[z^{-1}=\frac{1}{z}\] \[\frac{w}{z^{-2}}=wz^2\]

OpenStudy (unklerhaukus):

\[v^{1/2}=\sqrt{v}\]

OpenStudy (unklerhaukus):

\[\left(\frac{36a^{5/3}}{b^{2/5}c^{-1/3}}\right)^{1/2}\qquad\Longrightarrow\qquad \frac{\sqrt{36}a^{(5/3\times1/2)}}{b^{(2/5\times1/2)}c^{-1/3\times1/2}}\] \[\qquad\qquad\qquad\qquad\qquad\qquad\qquad\Downarrow\]